OpenStudy (anonymous):

The table below represents the displacement of a turtle from its nest as a function of time: Time (hours) x Displacement from nest (miles) y 0 5 1 27 2 49 3 71 4 93 Part A: What is the y-intercept of the function, and what does this tell you about the turtle? Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 4 hours, and tell what the average rate represents. Part C: What would be the domain of the function if the turtle continued to walk at this rate until it traveled 225 miles from the nest?

OpenStudy (anonymous):

I NEED HELP PLEASE ????????? HELP MEEEEEE

jimthompson5910 (jim_thompson5910):

`Part A: What is the y-intercept of the function, and what does this tell you about the turtle?` when x = 0, what is the value of y?

jimthompson5910 (jim_thompson5910):

The information is given in the table

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

do you see where I'm referring to?

OpenStudy (anonymous):

no sorry not that good at math

jimthompson5910 (jim_thompson5910):

look at the row that has x = 0 what is next to that value?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

so x = 0 and y = 5 pair up giving the ordered pair (0,5)

jimthompson5910 (jim_thompson5910):

what this means is that at time 0 hours, the distance is 5 miles basically the turtle is 5 miles away from the nest at the beginning

OpenStudy (anonymous):

ok so i write a equation

jimthompson5910 (jim_thompson5910):

no you just use words to explain what the y intercept means

jimthompson5910 (jim_thompson5910):

similar to how I explained it just now

OpenStudy (anonymous):

ok got it .......................so for part B I din't know how to do it can you help me

jimthompson5910 (jim_thompson5910):

`Part B: Calculate the average rate of change of the function represented by the table between x = 1 to x = 4 hours, and tell what the average rate represents.` fill in these blanks if x = 1, then y = ________ if x = 4, then y = ________

OpenStudy (anonymous):

Y=27 Y=93

jimthompson5910 (jim_thompson5910):

so we have these two points that are of interest to us (1,27) and (4,93)

jimthompson5910 (jim_thompson5910):

find the slope through those two points to get the answer

jimthompson5910 (jim_thompson5910):

slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

if thats not right then i think it might be m=-66/3

jimthompson5910 (jim_thompson5910):

you should get 66/3 which reduces to 22

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

what does this mean in terms of the problem?

OpenStudy (anonymous):

22 miles i think

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

not quite the full picture though

jimthompson5910 (jim_thompson5910):

slope = 22 = 22/1 this means that each time x increases by 1, y increases by 22 each time the hour amount goes up by 1, the mile amount goes up by 22 basically the turtle is going (22 miles)/(1 hour) = 22 mph. This is a very fast turtle lol

OpenStudy (anonymous):

yea lol thanks